OpenStudy (anonymous):

If an engineer were building a piece of machinery, why must she use a consistent system (either U.S. or metric) to measure all the parts of the machinery? Give a specific example of the errors that could occur if the engineer fails to do this.

OpenStudy (anonymous):

Theres one more after this one k.

OpenStudy (anonymous):

its clearly that u.s system involve:ft,lb,sec and si involve :meter,kilogram,s(second) knowing that 1ft=0.3048m and 1 lbf (force poun d)=4.44 N so. if we use metric system we should continue with metric and entierly we can change to us system by multiply in constant.

OpenStudy (anonymous):

thanks but can you help me wiith one more?

OpenStudy (anonymous):

Suppose a building inspector orders the height of a ceiling to be 8 feet tall, but the workers buy lumber that is only 50 centimeters tall. How many pieces of lumber will they have to stack to create an 8-foot ceiling? Explain your answer. (Hint: 1 inch = 2.54 centimeters.)

OpenStudy (anonymous):

8ft=8*12(in)=96in and 96*2.54=243 so we know that 8ft=243cm then =243/50=4.8 thats it

OpenStudy (anonymous):

woo your smart nice thanks !!!!!!!!!!!!!!!

OpenStudy (anonymous):

youre welcome